# arXiv:2603.01932v1[cs.CV] 02 Mar 2026

### v1

@@ 4 Results
We report results for the proposed radiance-index two-stream model on BAWSeg under three deployment-oriented protocols, namely within-plot, cross-plot, and cross-year. Performance is computed on held-out spatial blocks using mean Intersection over Union (mIoU), per-class IoU, micro-averaged precision, recall and F1, overall accuracy (OA), and Cohen’s $\kappa$ Cohen (1960). For mIoU confidence intervals (CI), we use a block bootstrap that resamples test blocks with replacement and recomputes the metric per replicate Efron and Tibshirani (1994). Unless stated otherwise, all metrics exclude ignore pixels.
@@ 4.1 Experimental Setting
Training configuration. Models were trained for 50 epochs on $256\times 256$ patches. Optimization used AdamW Loshchilov and Hutter (2019) with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.999$, weight decay $0.01$, and gradient clipping with a global norm of $1.0$. The initial learning rate was set to $6{\times}10^{-4}$ and decayed with a cosine schedule after a linear warm-up of 1500 iterations. The effective batch size was 16. Inputs were standardized per band or per index using training-set statistics. Data augmentation used spatial transforms only, including random horizontal and vertical flips and random rotations in $\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$, which preserves calibrated reflectance semantics.
Losses and inference. Training minimized a class-balanced cross-entropy on fused logits with an ignore mask, plus an auxiliary cross-entropy applied to the index stream with weight $\alpha{=}0.3$. Inference used single-scale evaluation at native patch resolution. For block-level evaluation, we applied a sliding-window over each held-out block using $256\times 256$ windows with stride 128 and averaged logits in overlapping regions before taking the argmax label map. This prevents boundary artifacts from dominating block metrics.
@@
Table 1: Within plot segmentation performance per year and field on held out test blocks. Mean IoU is estimated using $95\%$ bootstrap confidence intervals. Precision, recall and F1 are micro averaged over all classes.
@@
4.2 Segmentation Across Plots and Years
Table 1 reports within-plot performance stratified by year and field. Mean IoU remains in the range $[0.748,0.763]$ across the eight year-field combinations, with weed IoU consistently above 0.62. Micro-F1 remains close to 0.84, while OA remains close to 0.95 and $\kappa$ remains close to 0.79. These results indicate stable behavior across acquisition years for the same operational region and sensor configuration.
Table 2 summarizes protocol-level generalization. Under within-plot evaluation, the model reaches mIoU $0.756\pm 0.004$ with weed IoU 0.635. Under cross-plot transfer, training on one field and testing on the other yields mIoU near 0.71, and the primary degradation appears in the weed class, while crop IoU remains close to the within-plot level. Under cross-year testing, training on 2020–2022 and testing on 2023 yields mIoU $0.692\pm 0.007$ and weed IoU 0.544, with crop IoU remaining above 0.83. Across protocols, the relative sensitivity is concentrated in weed IoU, while OA stays above 0.93, reflecting the strong class imbalance between vegetation and the less frequent weed pixels.
@@ 4.2 Segmentation Across Plots and Years
Ours 2026 MSI $0.756\pm 0.004$ 0.847{0.847} 0.635{0.635} 0.794{0.794} 0.847{0.847} 0.946{0.946} 0.794{0.794} 22.8
4.3 Comparison with Existing Methods
Table 3 compares the proposed model with an index-based classical baseline and representative CNN and Transformer segmentation networks under the within-plot protocol. All deep baselines were trained and evaluated using the same block splits, patch size, optimizer type, learning-rate schedule family, and number of epochs as the proposed model. For multispectral inputs, we adapted the first layer of each architecture to accept five channels. For models that use patch embeddings, we replaced the input projection to accept five channels and initialized added channel weights by copying the mean of RGB weights when ImageNet pretraining was used, then fine-tuned end-to-end on BAWSeg.
@@ 4.3 Comparison with Existing Methods
Table 4: Ablation on the within‐plot protocol. $\Delta$ is relative to the full model. Params in millions. FLOPs in billions at $256\times 256$. Mem in GB at FP16, batch size 11. mIoU reports block-bootstrap 95% confidence intervals.
@@
w/o mean–slot broadcast $748\pm 0.007 0.0$08-0.008 0.622 12.3 32.4 2.50 80
@@
WSA heads h=4h{=}4 $750\pm 0.007 0.0$06-0.006 0.627 12.8 33.5 2.60 78
WSA heads h=8h{=}8 (full) $756\pm 0.004 0.0$000.000 0.635 12.8 33.6 2.60 78
WSA heads h=12h{=}12 $758\pm 0.007 +0.0$02+0.002 0.637 12.9 34.4 2.70 74
Mamba layers L=0L{=}0 $738\pm 0.007 0.0$18-0.018 0.610 12.2 31.5 2.50 82
Mamba layers L=1L{=}1 $746\pm 0.007 0.0$10-0.010 0.624 12.5 32.5 2.55 79
Mamba layers L=2L{=}2 (full) $756\pm 0.004 0.0$000.000 0.635 12.8 33.6 2.60 78
Mamba layers L=3L{=}3 $757\pm 0.007 +0.0$01+0.001 0.637 13.1 34.9 2.70 75
Notes. FLOPs: measured with fvcore (count MACs as FLOPs), input $1{\times}C{\times}256{\times}256$, eval() and Dropout off. Mem: PyTorch max-allocated GPU memory (FP16, batch=1=1). FPS: average over 200 iterations after 20 warm-up, timing wrapped by torch.cuda.synchronize(). Unless noted, differences $\leq\,$0.007 mIoU are not statistically significant under block-level bootstrap.
4.4 Ablation Study
Removing the vegetation-index modeling branch yields the largest accuracy drop, reducing mIoU from $0.756$ to $0.731$ and weed IoU from 0.635 to 0.596, while reducing parameters and increasing FPS. This isolates the contribution of index-based context beyond what is learned from calibrated multispectral bands alone. Within the index stream, removing relative positional bias in windowed self-attention reduces mIoU to $0.749$, and removing Slot Attention reduces mIoU to $0.744$ with a concurrent weed IoU decrease, indicating that grouping tokens into region descriptors contributes to weed discrimination under canopy mixing. Removing the mean-slot broadcast reduces mIoU to $0.748$, which shows that re-injecting a global descriptor into the spatial grid affects final predictions beyond the slot aggregation itself.
Varying attention heads shows a compute-accuracy trade-off. Reducing heads to h=4h{=}4 decreases mIoU to $0.750$ with a minor weed IoU decrease. Increasing heads to h=12h{=}12 yields mIoU $0.758$ while reducing FPS due to higher attention cost. For the state-space component, removing Mamba layers produces a clear reduction to $0.738$ mIoU and 0.610 weed IoU, and performance improves monotonically up to two layers, with limited change from two to three layers at increased compute and memory. Across these variants, improvements below the ablation CI scale in Table 4 should be interpreted as small under the block-bootstrap uncertainty, while changes above that scale correspond to consistent differences under the same evaluation procedure.
@@ 5 Discussion
The evaluation protocols expose two distinct regimes. When training and testing on disjoint spatial blocks from the same paddocks and seasons, VISA reaches 0.756 mIoU and 0.635 weed IoU, indicating that the model can separate weed patches from barley canopy at orthomosaic resolution while preserving row structure. Under cross-plot testing, mIoU drops to about 0.71 and the degradation concentrates on the weed class, while crop IoU remains close to the within-plot level. The same pattern holds under cross-year testing, where mIoU decreases to 0.692 and weed IoU to 0.544 with crop IoU still above 0.83. This behaviour is consistent with operational conditions in cereal systems, where crop appearance is relatively regular within a growth stage, while weed pressure, phenology, and within-row occlusion vary strongly across paddocks and seasons. The results therefore suggest that the remaining generalization gap is driven less by background and crop variation and more by shifts in weed morphology, density, and canopy mixing.
@@
A key design choice in VISA is to prevent the radiance cues and index cues from competing inside a single encoder. Vegetation indices suppress many illumination and exposure effects by construction and encode cross-band interactions that correlate with chlorophyll and canopy vigor, which makes them informative for vegetation delineation across flights. At the same time, indices compress radiometric detail and can attenuate subtle within-canopy contrasts that differentiate sparse weeds from barley leaves, especially when the canopy is dense and when soil is largely occluded. Conversely, calibrated multispectral reflectance retains high-frequency textures, row boundaries, and small spectral deviations that are informative for fine weed clusters, but this signal is more sensitive to residual radiometric drift and to local mosaic artefacts. VISA keeps these cues in separate pathways so that the radiance stream can preserve boundary detail without being dominated by normalized ratios, and the index stream can capture stable vegetation patterns without being distracted by band-specific noise. The fusion at native resolution then combines boundary precision from the radiance stream with context from the index stream, which aligns with the observed gains over a single-stream SegFormer baseline at similar scale.
@@
The ablation study helps interpret which components contribute to weed robustness. Removing the index branch reduces mIoU by 0.025 and lowers weed IoU from 0.635 to 0.596, indicating that vegetation-index reasoning contributes materially beyond what can be learned from calibrated bands alone. Within the index branch, windowed self-attention supports local structure and row-consistent patterns, and the relative positional bias yields a consistent mIoU reduction when removed, which is expected because row geometry is expressed primarily through relative offsets rather than absolute location in a tile. The state-space layers provide the largest single improvement among the index-branch modules, with a 0.018 mIoU drop when they are removed, which suggests that linear-time context propagation helps encode plot-scale regularities that extend beyond the attention window. Slot Attention further improves weed sensitivity, with a 0.012 mIoU drop when removed, consistent with the role of region-level grouping in stabilizing representations of sparse and fragmented weed patches embedded in crop canopy. The mean-slot broadcast contributes an additional margin, indicating that injecting a global descriptor back to the spatial grid helps resolve local ambiguity in mixed pixels.
@@
BAWSeg is constructed to make these conclusions meaningful under realistic deployment constraints. The dataset spans four seasons across two commercial paddocks, includes calibrated five-band orthomosaics and a consistent set of vegetation indices, and uses leakage-free spatial blocks for training and testing. Block splits are critical in orthomosaic learning because random sampling can place near-duplicate textures and lighting conditions in both training and test sets, yielding optimistic estimates that do not translate across paddocks. The cross-plot and cross-year protocols further stress the model under changes in weed distribution and season-specific appearance, which are the primary sources of failure in practice. The relatively stable crop IoU across protocols indicates that the calibration and preprocessing pipeline produces consistent crop appearance, while the weed IoU drop under shift highlights the need for modelling choices and training strategies that target rare, small, and occluded weed structures.
From a precision agriculture perspective, the reported performance is most relevant when translated into actionable maps. The model outputs per-pixel posteriors that can be aggregated into patch-level weed objects and converted into prescription maps using minimum-area constraints that match sprayer resolution and drift tolerance. In this setting, false positives are often more costly than small boundary errors, because over-spraying increases chemical use and can accelerate resistance. The confidence maps produced by the softmax output enable risk-aware filtering, where low-confidence weed detections can be flagged for targeted scouting rather than immediate treatment. The inference throughput reported in the ablation table indicates that the architecture is compatible with practical offline processing of field-scale orthomosaics on a single GPU, which matches common agronomic workflows that generate maps after flight rather than on the UAV.
@@
Several limitations remain. First, the geographic scope is limited to two paddocks in one grainbelt region and one UAV sensor configuration, so the reported cross-year robustness does not fully characterize transfer to different soil types, management practices, or imaging systems. Second, the labels represent three classes and resolve crop and weed overlap by polygon precedence, which is operationally convenient but does not capture the fractional composition of mixed pixels in dense canopy. This affects both training targets and evaluation, especially when weeds are small and partially occluded. Third, orthomosaic generation can introduce seamline artefacts and local blur that change with overlap and wind conditions, which can act as nuisance variation across campaigns. Addressing these constraints motivates several concrete directions. Expanding BAWSeg to additional farms and sensors would improve coverage of radiometric and agronomic variability. Introducing probabilistic or soft labels near ambiguous boundaries and in mixed canopy regions would better match the underlying physical mixture. For generalization, domain generalization and test-time adaptation strategies that operate on calibrated reflectance and indices are promising, since the cross-plot and cross-year drops are concentrated in the weed class. Finally, weakly supervised extensions that exploit unlabeled mosaics from new seasons could reduce annotation burden while improving robustness to the shifts that are most salient in deployment.

### v2

@@ 4 Results
We report results for the proposed radiance-index two-stream model on BAWSeg under three deployment-oriented protocols, namely within-plot, cross-plot, and cross-year. Performance is computed on held-out spatial blocks using mean Intersection over Union (mIoU), per-class IoU, micro-averaged precision, recall and F1, overall accuracy (OA), and Cohen’s $\kappa$ Cohen (1960). For mIoU confidence intervals (CIs), we use a block bootstrap that resamples test blocks with replacement and recomputes the metric per replicate Efron and Tibshirani (1994). Unless stated otherwise, all metrics exclude ignore pixels.
@@ 4.1 Experimental Setting
Training configuration. Models were trained for 50 epochs on $256\times 256$ patches. Optimization used AdamW Loshchilov and Hutter (2019) with $\beta_{1}~{=}~0.9$, $\beta_{2}~{=}~0.999$, weight decay $0.01$, and gradient clipping with a global norm of $1.0$. The initial learning rate was set to $6~{\times}~10^{-4}$ and decayed with a cosine schedule after a linear warm-up of 1500 iterations. The effective batch size was 16. Inputs were standardized per band or per index using training-set statistics. Data augmentation used spatial transforms only, including random horizontal and vertical flips and random rotations in $\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$, which preserves calibrated reflectance semantics.
Losses and inference. Training minimized a class-balanced cross-entropy on fused logits with an ignore mask, plus an auxiliary cross-entropy applied to the index stream with weight $\alpha~{=}~0.3$. Inference used single-scale evaluation at native patch resolution. For block-level evaluation, we applied a sliding window over each held-out block using $256\times 256$ windows with stride 128 and averaged logits in overlapping regions before taking the argmax label map. This prevents boundary artifacts from dominating block metrics.
4.2 Segmentation Across Plots and Years
@@
Table 1 reports within-plot performance stratified by year and field. Mean IoU remains in the range $[0.748,0.763]$ across the eight year–field combinations, with weed IoU consistently above 0.62. Micro-F1 remains close to 0.84, while OA remains close to 0.95 and $\kappa$ remains close to 0.79. These results indicate stable behavior across acquisition years for the same operational region and sensor configuration.
Table 1: Within-plot segmentation performance per year and field on held-out test blocks. Mean IoU is estimated using $95\%$ bootstrap confidence intervals. Precision, recall and F1 are micro-averaged over all classes.
@@
Table 2 summarizes protocol-level generalization. Under within-plot evaluation, the model reaches mIoU $0.756\pm 0.004$ with weed IoU 0.635. Under cross-plot transfer, training on one field and testing on the other yields mIoU near 0.71, and the primary degradation appears in the weed class, while crop IoU remains close to the within-plot level. Under cross-year testing, training on 2020–2022 and testing on 2023 yields mIoU $0.692\pm 0.007$ and weed IoU 0.544, with crop IoU remaining above 0.83. The larger drop in the cross-year setting is mainly attributable to temporal domain shift across acquisition years, which changes weed morphology, weed density, canopy mixing, exposed-soil proportion, and residual illumination conditions in the orthomosaics. This effect is concentrated in the weed class, while crop IoU remains comparatively stable, indicating that the degradation is driven primarily by year-dependent variation in weed appearance rather than by a general failure of preprocessing or radiometric calibration. Across protocols, the relative sensitivity is concentrated in weed IoU, while OA stays above 0.93, reflecting the strong class imbalance between vegetation and the less frequent weed pixels.
@@
4.3 Comparison with Existing Methods
Table 3 compares the proposed model with an index-based classical baseline and representative CNN and Transformer segmentation networks under the within-plot protocol. All deep baselines were trained and evaluated using the same block splits, patch size, optimizer type, learning-rate schedule family, and number of epochs as the proposed model. For multispectral inputs, we adapted the first layer of each architecture to accept five channels. For models that use patch embeddings, we replaced the input projection to accept five channels and initialized added channel weights by copying the mean of RGB weights when ImageNet pretraining was used, and then fine-tuned end-to-end on BAWSeg.
@@ 4.3 Comparison with Existing Methods
Ours 2026 MSI $0.756\pm 0.004$ 0.847 0.635 0.794 0.847 0.946 0.794 22.8
@@
4.4 Ablation Study
Table 4: Ablation on the within‐plot protocol. $\Delta$ is relative to the full model. Params in millions. FLOPs in billions at $256\times 256$. Mem in GB at FP16, batch size 11. mIoU reports block-bootstrap 95% confidence intervals. Arrows indicate whether higher or lower values are preferable.
@@ 4.4 Ablation Study
w/o mean-slot broadcast $748\pm 0.007 0.0$08-0.008 0.622 12.3 32.4 2.50 80
@@
WSA heads h=4h~{=}~4 $750\pm 0.007 0.0$06-0.006 0.627 12.8 33.5 2.60 78
WSA heads h=8h~{=}~8 (full) $756\pm 0.004 0.0$000.000 0.635 12.8 33.6 2.60 78
WSA heads h=12h~{=}~12 $758\pm 0.007 +0.0$02+0.002 0.637 12.9 34.4 2.70 74
Mamba layers L=0L~{=}~0 $738\pm 0.007 0.0$18-0.018 0.610 12.2 31.5 2.50 82
Mamba layers L=1L~{=}~1 $746\pm 0.007 0.0$10-0.010 0.624 12.5 32.5 2.55 79
Mamba layers L=2L~{=}~2 (full) $756\pm 0.004 0.0$000.000 0.635 12.8 33.6 2.60 78
Mamba layers L=3L~{=}~3 $757\pm 0.007 +0.0$01+0.001 0.637 13.1 34.9 2.70 75
Notes. FLOPs: measured with fvcore (count MACs as FLOPs), input $1~{\times}~C~{\times}~256~{\times}~256$, eval() and Dropout off. Mem: PyTorch max-allocated GPU memory (FP16, batch=1=1). FPS: average over 200 iterations after 20 warm-ups, timing wrapped by torch.cuda.synchronize(). Unless noted, differences $\leq\,$0.007 mIoU are not statistically significant under block-level bootstrap.
Removing the vegetation-index modelling branch yields the largest accuracy drop, reducing mIoU from $0.756$ to $0.731$ and weed IoU from 0.635 to 0.596, while reducing parameters and increasing FPS. This isolates the contribution of index-based context beyond what is learned from calibrated multispectral bands alone. Within the index stream, removing relative positional bias in windowed self-attention reduces mIoU to $0.749$, and removing Slot Attention reduces mIoU to $0.744$ with a concurrent weed IoU decrease, indicating that grouping tokens into region descriptors contributes to weed discrimination under canopy mixing. Removing the mean-slot broadcast reduces mIoU to $0.748$, which shows that re-injecting a global descriptor into the spatial grid affects final predictions beyond the slot aggregation itself.
Varying attention heads shows a compute–accuracy trade-off. Reducing heads to h=4h~{=}~4 decreases mIoU to $0.750$ with a minor weed IoU decrease. Increasing heads to h=12h~{=}~12 yields mIoU $0.758$ while reducing FPS due to higher attention cost. For the state-space component, removing Mamba layers produces a clear reduction to $0.738$ mIoU and 0.610 weed IoU, and performance improves monotonically up to two layers, with limited change from two to three layers at increased compute and memory. Across these variants, improvements below the ablation CI scale in Table 4 should be interpreted as small under the block-bootstrap uncertainty, while changes above that scale correspond to consistent differences under the same evaluation procedure.
@@ 5 Discussion
The evaluation protocols expose two distinct regimes. When training and testing on disjoint spatial blocks from the same paddocks and seasons, VISA reaches 0.756 mIoU and 0.635 weed IoU, indicating that the model can separate weed patches from barley canopy at orthomosaic resolution while preserving row structure. Under cross-plot testing, mIoU drops to about 0.71 and the degradation concentrates on the weed class, while crop IoU remains close to the within-plot level. The same pattern holds under cross-year testing, where mIoU decreases to 0.692 and weed IoU to 0.544, with crop IoU still above 0.83. This behaviour is consistent with operational conditions in cereal systems, where crop appearance is relatively regular within a growth stage, while weed pressure, phenology, and within-row occlusion vary strongly across paddocks and seasons. Therefore, the results suggest that the remaining generalization gap is driven less by background and crop variation and more by shifts in weed morphology, density, and canopy mixing.
@@
A key design choice in VISA is to prevent the radiance cues and index cues from competing inside a single encoder. Vegetation indices suppress many illumination and exposure effects by construction and encode cross-band interactions that correlate with chlorophyll and canopy vigor, which makes them informative for vegetation delineation across flights. At the same time, indices compress radiometric detail and can attenuate subtle within-canopy contrasts that differentiate sparse weeds from barley leaves, especially when the canopy is dense and when soil is largely occluded. Conversely, calibrated multispectral reflectance retains high-frequency textures, row boundaries, and small spectral deviations that are informative for fine weed clusters, but this signal is more sensitive to residual radiometric drift and to local mosaic artefacts. VISA keeps these cues in separate pathways so that the radiance stream can preserve boundary detail without being dominated by normalized ratios, and the index stream can capture stable vegetation patterns without being distracted by band-specific noise. The fusion at native resolution then combines boundary precision from the radiance stream with context from the index stream, which aligns with the observed gains over a single-stream SegFormer baseline at a similar scale.
@@
The ablation study helps interpret which components contribute to weed robustness. Removing the index branch reduces mIoU by 0.025 and lowers weed IoU from 0.635 to 0.596, indicating that vegetation-index reasoning contributes materially beyond what can be learned from calibrated bands alone. Within the index branch, windowed self-attention supports local structure and row-consistent patterns, and the relative positional bias yields a consistent mIoU reduction when removed, which is expected because row geometry is expressed primarily through relative offsets rather than absolute location in a tile. The state-space layers provide the largest single improvement among the index-branch modules, with a 0.018 mIoU drop when they are removed, which suggests that linear-time context propagation helps encode plot-scale regularities that extend beyond the attention window. Slot Attention further improves weed sensitivity, with a 0.012 mIoU drop when removed, consistent with the role of region-level grouping in stabilizing representations of sparse and fragmented weed patches embedded in crop canopies. The mean-slot broadcast contributes an additional margin, indicating that injecting a global descriptor back to the spatial grid helps resolve local ambiguity in mixed pixels.
The sequential component in the index branch also merits clarification with respect to scan direction. In the current model, the Mamba-style state-space blocks use a single row-major raster scan to convert the 2D feature grid into a 1D sequence. We do not use bidirectional or cross-scanning in this study. A multi-directional design could, in principle, reduce directional bias further for anisotropic field structure. However, in the present architecture, the sequential module is preceded by windowed self-attention with relative positional bias, and the training pipeline includes rotation and flip augmentation, which already reduce sensitivity to a fixed scan order. In addition, the inputs are orthorectified and radiometrically calibrated before sequence construction, so residual viewing and illumination effects are primarily handled in preprocessing rather than by changing scan direction. Under this setting, we did not observe evidence in the protocol-level results that scan order was a dominant source of error. Therefore, we retained the single-scan formulation to preserve the lightweight design of the index branch, while regarding multi-directional scanning as a possible extension for future work.
The selected index set should also be distinguished from the vegetation-index products provided in standard DJI software workflows. For Phantom 4 Multispectral data, DJI Terra provides NDVI, GNDVI, NDRE, LCI, and OSAVI, which only partially overlap with the index set used in this study. Our choice of NDVI, GNDVI, EVI, SAVI, and MSAVI was guided by the need to retain one blue-band index and two soil-adjusted indices that are directly relevant to mixed-canopy and exposed-soil conditions in BAWSeg. In future deployment-oriented settings, vendor-generated indices could be used as an alternative index source if one aims to simplify the preprocessing workflow. However, they would define a different input space from the one used here and, therefore, should be treated as a new variant that requires retraining and revalidation rather than as a direct replacement.
The platform also provides a co-registered RGB stream, which could be combined with multispectral data in future extensions. In such a setting, RGB would mainly contribute visible texture, edge, and morphological detail, while the multispectral bands provide red edge and near-infrared responses that are more informative for vegetation condition and crop–weed discrimination under mixed canopy and illumination variation. Their combination is, therefore, most relevant for boundary refinement, delineation of small weed patches, and visually ambiguous regions where appearance and spectral cues are both needed. We did not include RGB fusion in the present study because our objective was to establish a controlled benchmark for multispectral reflectance and vegetation-index modelling only. Adding RGB would define a different multimodal setting and should be evaluated as a separate extension under the same block-based protocols.
@@
BAWSeg is constructed to make these conclusions meaningful under realistic deployment constraints. The dataset spans four seasons across two commercial paddocks, includes calibrated five-band orthomosaics and a consistent set of vegetation indices, and uses leakage-free spatial blocks for training and testing. Block splits are critical in orthomosaic learning because random sampling can place near-duplicate textures and lighting conditions in both training and test sets, yielding optimistic estimates that do not translate across paddocks. The cross-plot and cross-year protocols further stress the model under changes in weed distribution and season-specific appearance, which are the primary sources of failure in practice. The relatively stable crop IoU across protocols indicates that the calibration and preprocessing pipeline produces consistent crop appearance, while the weed IoU drop under shift highlights the need for modelling choices and training strategies that target rare, small, and occluded weed structures. This limitation is related to both acquisition scale and preprocessing. At the nominal multispectral GSD of about 6.35cm/pixel6.35\ $\text${cm/pixel} at 120 m altitude, BAWSeg is well suited to row-level and patch-level weed segmentation, but very small weeds near the spatial resolution limit are inherently more difficult to separate from crop canopy and mixed pixels. The $3\times 3$ median filter used in preprocessing improves robustness to isolated pixel noise and helps stabilize subsequent alignment and reflectance processing, but it can also attenuate one to two pixel weed responses before feature extraction. Therefore, we regard the current preprocessing as a practical trade-off that improves overall orthomosaic quality while leaving very small weed detectability as a residual limitation of the pipeline.
From a precision agriculture perspective, the reported performance should be interpreted together with the efficiency results in Table 4. The full model uses 22.8 M parameters, 33.6 GFLOPs, 2.60 GB FP16 memory, and runs at 78 FPS for $256\times 256$ patches on a single RTX 4090. These results indicate that the model is practical for routine post-flight deployment on a standard workstation, where tiled orthomosaics can be segmented efficiently for same-cycle weed mapping, targeted scouting, and prescription preparation. In this workflow, the dominant time cost is more likely to come from orthomosaic generation and preprocessing rather than from network inference itself. The confidence maps produced by the softmax output further support risk-aware filtering, where low-confidence weed detections can be flagged for manual inspection instead of immediate treatment. This makes the current system more suitable for reliable offline decision support than for direct onboard UAV inference, which is consistent with common agronomic practice.
@@
Several limitations remain. First, the geographic scope is limited to two paddocks in one grainbelt region and one UAV sensor configuration, so the reported cross-year robustness does not fully characterize transfer to different soil types, management practices, or imaging systems. Second, the labels represent three classes and resolve crop and weed overlap by polygon precedence, which is operationally convenient but does not capture the fractional composition of mixed pixels in dense canopy. This affects both training targets and evaluation, especially when weeds are small and partially occluded. Third, orthomosaic generation can introduce seamline artefacts and local blur that change with overlap and wind conditions, which can act as nuisance variation across campaigns. Addressing these constraints motivates several concrete directions. Expanding BAWSeg to additional farms and sensors would improve coverage of radiometric and agronomic variability. Introducing probabilistic or soft labels near ambiguous boundaries and in mixed-canopy regions would better match the underlying physical mixture. For generalization, domain generalization and test-time adaptation strategies that operate on calibrated reflectance and indices are promising, since the cross-plot and cross-year drops are concentrated in the weed class. Finally, weakly supervised extensions that exploit unlabeled mosaics from new seasons could reduce annotation burden while improving robustness to the shifts that are most salient in deployment.